OpenStudy (anonymous):

Karina uses the system of equations below to compare the monthly utility costs in July and December for electricity, x, and natural gas, y. 750x + 17y = 141.61 300x + 30y = 75.90 Karina solves the system using linear combination and arrives at the equation 116y = 96.28. She then solves this equation for y. Which statement explains Karina’s solution? a. The cost of electricity is $0.17 per unit. b. The cost of natural gas is $0.20 per unit. c. The cost of electricity is $0.72 per unit. d. The cost of natural gas is $0.83 per unit.

OpenStudy (anonymous):

@cj49

OpenStudy (cj49):

116y=96.28 y=96.28/116=0.83

OpenStudy (cj49):

wait... multiply the first equation by 30 and the second by 17 and then subtract 1 from 2 22500x+510y=4248.3 51500x+510y=2958 ull gt x=0.17 and y=0.83